OpenStudy (anonymous):

(64)^1/3 X (128)^2/7=

OpenStudy (aum):

\[\large 64 = 2^6 \\ \large 128 = 2^7 \\ \large (64)^{1/3} * (128)^{2/7} = (2^6)^{1/3} * (2^7)^{2/7} = 2^2 * 2^2 = 2^{2+2} = 2^4 = 16 \]

OpenStudy (anonymous):

ooooooooooooh okay. thank you